# LES-Talker: Fine-Grained Emotion Editing for Talking Head Generation in Linear Emotion Space

### v1

@@ Abstract
11footnotetext: These authors contributed equally.22footnotetext: * Corresponding author.
@@ 1 Introduction
Talking head generation has attracted significant attention from researchers in recent years and has many applications in the fields of digital human creation [45], virtual reality [7], etc. To enhance the diversity and expressiveness of talking head generation, emotion editing tasks were emphasized. However, existing studies have the following two notable drawbacks: (1) the interpretability of methods is lacking; (2) the effectiveness of fine-grained emotion editing is still unsatisfactory. On the one hand, some studies offer expression reference images without clear emotion definition [33, 34, 20], while others employ discrete emotion labels [17, 32, 24]. Additionally, some extract emotion features from latent spaces [34, 25] leading to an implicit transformation of emotion. Although Facial Action Units (AUs) effectively describe emotions, various methods [13, 16, 14] use different combinations of AUs for the same emotion. Moreover, AUs alone remain insufficient. On the other hand, several studies [35, 25, 33] focus on generating videos with specific emotions, realizing only coarse-grained emotion editing. One of the few fine-grained emotion studies [34, 31] also shows only discrete emotion levels in three emotion types.
@@ 2.1 Talking Head Video Generation
Existing research on talking head generation is mainly categorized into audio-driven [36, 41, 30, 1, 19, 6] and video-driven [23, 37, 12, 46] two broad categories. In audio-driven methods, lightweight use of video to compensate the lack of information is a common approach, such as StyleTalk [24] extracting facial motion patterns using a transformer, and PC-AVS [48] incorporating head pose information through implicit low-dimensional pose coding. Video-driven methods utilize the richer information contained in the input video to generate more natural results, focusing on different driving methods and intermediate representations, which are divided into implicit and explicit. Implicit modeling represents scenes by learning mathematical functions, such as Signed Distance Functions (SDF) [38, 47, 29] and Neural Radiation Fields (NeRF) [18, 21, 28]. Explicit modeling directly constructs editable 3D geometric representations, such as 3D Morphable Model (3DMM) [42, 27, 39] and 3D Gaussian Splatting (3D-GS) [40, 9, 22].
@@ 2.2 Fine-grained Facial Emotion Video Generation
In talking head generation, early studies [26, 8] focused on improving visual quality, while recent efforts have explored coarse-grained facial emotion editing. EmoTalk [25] employs an audio-driven method using two different audio extractors to separate emotion and content, and EDTalk [33] emphasizes controlled talking head generation by decoupling mouth shape, head pose, and emotional expression. However, these studies only achieve coarse-grained emotion expression.Recently, a few fine-grained emotion studies have emerged. Under co-driven conditions, PD-FGC [34] demonstrates linear interpolation between expression features from different sources by learning the expression latent space. EmoSpeaker [15] decouples audio input into content and emotion vectors, editing content vectors for different emotions. FG-EmoTalk [31] shows video-driven realization of three defined emotions and several levels of fine-grained emotions using AUs as control inputs. However, these studies lack clarity on the principles of fine-grained emotion changes and primarily realize discrete rather than continuous changes.
@@ 4.2 Editing Effectiveness
MEAD 146.454 0.469 14.864 0.191 11.957 2.674 -2.000 2.73 3.42 3.87 3.72
EVP 56.650 0.453 16.308 0.341 12.443 3.163 5.000 2.83 3.55 3.71 3.85
EAMM 204.002 0.396 12.832 0.135 10.091 3.046 -4.000 3.12 3.72 3.65 3.56
SadTalker 25.746 0.718 20.728 0.249 8.875 6.484 1.000 3.71 - 3.30 3.86
@@ 4.3 Visual Quaility
Comparision with Other Methods. To the best of our knowledge, no other research in talking head generation explicitly defines fine-grained emotions. We focus on visual quality, which is tested randomly in the neutral generation case of the MEAD dataset. Merely generating lip movements is insufficient for real-world applications. We compare our method with state-of-the-art emotion-driven facial expression generation methods, such as EAMM, EVP, and MEAD, as well as the one-shot generation method SadTalker. Our method demonstrates improved visual quality in Fig. 10. Tab. 1 shows our method exhibits superior performance in video quality and lip synchronization.
@@
User Study. We conducted user studies with 15 participants to evaluate our model across multiple metrics (rated 1–5) compared to other methods, as detailed in Tab.1. Our method outperformed others. As illustrated in Tab.2, we assessed user-perceived emotion levels when presented with videos labeled $(emo,level)$.

### v2

@@ 1 Introduction
Talking head generation has attracted significant attention from researchers in recent years and has many applications in the fields of digital human creation [51], virtual reality [8], etc. To enhance the diversity and expressiveness of talking head generation, emotion editing tasks were emphasized. However, existing studies have the following two notable drawbacks: (1) the interpretability of methods is lacking; (2) the effectiveness of emotion editing is still remains at a coarse-grained level. On the one hand, some studies offer expression reference images without clear emotion definition [37, 38, 22], while others employ discrete emotion labels [18, 36, 28]. Additionally, some extract emotion features from latent spaces [38, 29, 23] leading to an implicit transformation of emotion. Although Facial Action Units (AUs) effectively describe emotions, various methods [14, 17, 15] use different combinations of AUs for the same emotion. Moreover, AUs alone remain insufficient. On the other hand, several studies [40, 29, 37] focus on generating videos with specific emotions, realizing only coarse-grained emotion editing. One of the few fine-grained emotion studies [38, 35] also shows only discrete emotion levels in three emotion types.
@@ 2.1 Talking Head Video Generation
Existing research on talking head generation is mainly categorized into audio-driven [41, 46, 34, 1, 20, 7] and video-driven [27, 42, 13, 52] two broad categories. In audio-driven methods, lightweight use of video to compensate the lack of information is a common approach, such as StyleTalk [28] extracting facial motion patterns using a transformer, and PC-AVS [54] incorporating head pose information through implicit low-dimensional pose coding. Video-driven methods utilize the richer information contained in the input video to generate more natural results, focusing on different driving methods and intermediate representations, which are divided into implicit and explicit. Implicit modeling represents scenes by learning mathematical functions, such as Signed Distance Functions (SDF) [43, 53, 33] and Neural Radiation Fields (NeRF) [19, 24, 32]. Explicit modeling directly constructs editable 3D geometric representations, such as 3D Morphable Model (3DMM) [48, 31, 44] and 3D Gaussian Splatting (3D-GS) [45, 10, 25, 6].
@@ 2.2 Fine-grained Facial Emotion Video Generation
In talking head generation, early studies [30, 9] focused on improving visual quality, while recent studies [47, 39, 26] have explored coarse-grained facial emotion editing. EmoTalk [29] employs an audio-driven method using two different audio extractors to separate emotion and content, and EDTalk [37] emphasizes controlled talking head generation by decoupling mouth shape, head pose, and emotional expression. However, these studies only achieve coarse-grained emotion expression.Recently, a few fine-grained emotion studies have emerged. Under co-driven conditions, PD-FGC [38] demonstrates linear interpolation between expression features from different sources by learning the expression latent space. EmoSpeaker [16] decouples audio input into content and emotion vectors, editing content vectors for different emotions. FG-EmoTalk [35] shows video-driven realization of three defined emotions and several levels of fine-grained emotions using AUs as control inputs. However, these studies lack clarity on the principles of fine-grained emotion changes and primarily realize discrete rather than continuous changes.
@@ 4.2 Editing Effectiveness
MEAD[40] 146.454 0.469 14.864 0.191 11.957 2.674 -2.000 2.73 3.42 3.87 3.72
EVP[21] 56.650 0.453 16.308 0.341 12.443 3.163 5.000 2.83 3.55 3.71 3.85
EAMM[22] 204.002 0.396 12.832 0.135 10.091 3.046 -4.000 3.12 3.72 3.65 3.56
SadTalker[48] 25.746 0.718 20.728 0.249 8.875 6.484 1.000 3.71 - 3.30 3.86
EAT[18] 135.439 0.706 13.139 0.228 7.531 8.164 -2.000 3.38 3.76 3.91 3.93
@@ 4.3 Visual Quaility
We compare our method with state-of-the-art open-source emotion-driven models (EAMM, EVP, EAT) and the one-shot method SadTalker, achieving superior visual quality.
@@
User Study. We conducted user studies with 15 participants to evaluate our model across multiple metrics (rated 1–5) compared to other methods, as detailed in Tab.1. Our method outperformed others. As illustrated in Tab.2, we assessed user-perceived emotion levels when presented with videos labeled $(emo,level)$. The results confirm a consistent upward trend between theoretical and perceived emotion intensity, validating our framework’s effectiveness. Meanwhile, intensity shifts in sadness and surprise are subtler than in happiness and anger.
@@ 5 Conclusion
Appendix
Appendix A Further Description of LES
AUs Action Description AUs Action Description
AU1 Inner Brow Raiser AU14 Dimpler
AU2 Outer Brow Raiser AU15 Lip Corner Depressor
AU4 Brow Lowerer AU17 Chin Raiser
AU5 Upper Lid Raiser AU20 Lip Stretcher
AU6 Cheek Raiser AU23 Lip Tightener
AU7 Lid Tightener AU25 Lips Part
AU9 Nose Wrinkler AU26 Jaw Drop
AU10 Upper Lip Raiser AU45 Blink
AU12 Lip Corner Puller
Table 4: Facial Action Units Description.
A.1 Physical Meaning
Our proposed Linear Emotion Space contains a total of 41 dimensions. The first 17 dimensions, which make up the Action Subspace $\mathbb{A}$, directly correspond to the 17 Facial Action Units (AUs) listed above. Their values have a unified physical meaning across all emotions, representing the amplitude of movement of each facial unit. The next 17 dimensions (18-34) form the Isolation Subspace $\mathbb{I}$ and also correspond to the same 17 AUs. However, in this subspace, the values have distinct physical meanings under each emotion, representing the magnitude of fluctuation in the movement of each facial unit. Finally, the last 7 dimensions (35-41) of the Isolation Subspace are dedicated to isolating individual emotions and directly capturing the overall intensity of each emotion.
Cluster Top Emotion-Level Pairs Percentage
0 happy_level_3 (46.67%) 4.55
1 contempt_level_3 (32.40%) 10.81
2 disgusted_level_3 (42.24%) 7.81
3 fear_level_3 (39.32%) 7.88
4 angry_level_3 (25.85%) 22.66
5 neutral_level_1 (48.41%) 27.54
6 sad_level_3 (37.66%) 10.37
7 happy_level_3 (68.67%) 8.38
Table 5: GMM Clusters for vectors in ACT Subspace.
Cluster Top Emotion-Level Pairs Percentage
0 neutral-level_1 (42.63%) 33.80
1 sad-level_3 (38.66%) 13.06
2 happy-level_3 (59.41%) 3.40
3 sad-level_3 (23.49%) 5.59
4 contempt-level_3 (27.73%) 12.63
5 contempt-level_3 (40.43%) 12.66
6 disgusted-level_3 (42.16%) 9.66
7 happy-level_3 (64.10%) 9.19
Table 6: KMeans Clusters for vectors in ACT Subspace.
A.2 Cluster Analysis
As shown in Tab. 5 and Tab. 6, we provide a detailed demonstration of the clustering analysis mentioned in the main text. When performing emotion reconstruction using only the Action Subspace, both K-means and GMM clustering exhibit poor results. Within each cluster, the most frequent emotion-level pair does not significantly dominate. Across clusters, the sample points of each class—which should be approximately equal in number—are unevenly distributed, and the different emotions cannot be effectively separated.
Figure 12: The complete outlier matrix.
A.3 Complete Outlier Matrix
As illustrated in Fig. 12, we present the complete outlier matrix. The horizontal axis corresponds to the indices of the 17 controllable action facial units, while the vertical axis denotes the emotion-level pairs. The outlier matrix is precomputed prior to training using the LES definitions described in the main text, leveraging a pre-trained model and statistical methods. The matrix provides 22 feature vectors $\mathbf{uf}$ in the Action Subspace, which are employed for emotion transformation in the Emotion Injector.
Figure 13: Unbalanced distribution before standardization.
Figure 14: Od distribution in ISO Subspace.
A.4 Distribution Optimizing
As shown in Fig. 13, prior to applying the Opt1 and Opt2 methods defined in the LES, all AUs exhibit an unbalanced distribution. This is partly due to inherent limitations of the dataset itself, and partly due to defects in the AU extraction tools. Such an unbalanced distribution leads to the inability to establish a unified representation in the subsequent LES-defined space, and also results in deficiencies in the final visual effects. As illustrated in Fig. 14, after distribution optimization, in the Isolation Space, the vector distance $Od$ used to represent emotional levels achieves a uniform distribution across different emotions, providing support for a unified level representation of different emotions.
Appendix B Training Strategy
B.1 Coarse-to-Fine Strategy
During training, a two-step strategy from coarse to fine was employed as follows:
Coarse Training: The coarse training aligns with the design concept of the Action Subspace, maximizing the use of AUs for facial action representation. Train the first-level CDAN individually using the first frame of each video as the reference image with $\boldsymbol{u}\in\mathbb{A}$. Since the first frame better matches the identity information of subsequent frames, this training approach provides more precise facial action control.
Fine Training: As using only AUs is insufficient, facial expressions beyond AUs need to be further adjusted through the Isolation Subspace in the fine training. Freeze the first CDAN, then train the second-level CDAN using the neutral emotion frame as the reference image with $\boldsymbol{v}\in\mathbb{I}$. Because the final goal is to perform arbitrary emotion transformations starting from a neutral emotion, our method enables the network to have stronger emotion transformation capabilities.
Finally, freeze both levels of CDAN and train the Offset Decoder separately, utilizing information from the audio modality to compensate for the residuals between the remaining predicted coefficients and the ground truth coefficients. This specialized training method not only allows the model to achieve the desired emotional expression capabilities but also attains higher visual quality.
B.2 Hyper-parameters
To ensure rapid and stable convergence of the model training towards lower loss values, we employed the Adam optimizer and conducted extensive hyperparameter tuning, including adjustments to the factor and patience parameters. Experimental results demonstrated that the initial learning rate for each epoch was set as $lr\times decay^{\text{epoch}}$, where $lr\in\{10^{-3},\mathbf{10}^{-4},10^{-5}\}$ and $decay\in\{0.9,\mathbf{0.86},0.74\}$. The batch size was configured as $\{\mathbf{10},20,30\}$, and patience was set to $\{1500,\mathbf{15000},30000\}$. Larger batch sizes and excessively low learning rates resulted in prolonged training times, while larger factors and learning rates compromised training stability. Additionally, smaller decay rates and patience values led to insufficient convergence. Training was terminated after 30 epochs.
      Guide Space       Method       FLD↓
      ACT       MLP       1.959
      ACT       CDAN       1.150
      ISO       MLP       2.230
      ISO       CDAN w/o serial       2.012
      ISO       CDAN       1.421
Table 7: Facial Landmark Distance (FLD) under emotional conditions, based on 3DMM, summed over 200 frames.
Appendix C Further Experiment
C.1 Facial Landmark Distance
Figure 15: Generation in ablation study.
Figure 16: Multi-perspective visualization of the generated results using t-SNE downscaling.
As shown in Tab. 7, we utilized the Facial Landmark Distance (FLD) based on 3DMM coefficients, randomly selecting 200 frames for evaluation. Initially, we tested the Action Subspace independently, using reference images with identical emotions and identities for generation. Subsequently, we incorporated the Isolation Subspace, employing reference images with varying emotions and identities for generation. Replacing CDAN with a three-layer multilayer perceptron (MLP) or removing the serial connections resulted in a significant difference in FLD.
C.2 Visual Emotion Effects
Fig. 15 illustrates the more detailed generative effects of LES-Talker compared to ablation results. Surprised* and angry* indicate the ablation where both levels of CDAN are replaced with a 3-layer MLP (consistent with Tab.7) . Sad’, Sad” and Sad”’ respectively, indicate the absence of a serial connection between CDANs, the second level CDAN, and the special training strategy. Emotion level is set to 3.00. When either the CDANs, the special training strategy, or the serial connection is removed, the model loses the ability to express distinct emotional details, causing angry, surprised, and sad to resemble happy-like emotions, primarily by widening the mouth as the emotion level increases.
C.3 Visualization of Emotion Representations
We generated 1,755 videos across five emotions, with levels ranging from 0 to 4 in increments of 0.0114. Each video’s frames were converted into 41-dimensional LES representations, averaged, and reduced using t-SNE. Fig. 16 displays the multi-view spatial visualizations. For any pair of emotional point clouds, a separating plane exists in at least one view. Within each point cloud, intensity variations are clear and uniform. Our LES design achieves superior emotional isolation and fine-grained representation.
Appendix D Core Inference Code
D.1 Emotion Injector
We present the inference process of the Emotion Injector, which primarily relies on the emotional semantics inherent in each vector within the designed LES space. The Emotion Injector utilizes the LES representations generated for each frame, along with 22 feature vectors stored in the Action Subspace corresponding to eight emotions (including neutral), to perform emotion transformations. In the absence of an AU Source, a preset neutral emotion is employed as the initial LES representation.
Algorithm 1 Emo Injector Algorithm
Input: emo_intensity, emo_name, mode_type, emo_control_pre
Output: emo_control, point_control
1:  Load stats from CSV
2:  Initialize emo_control_dict and point_control_dict
3:  if emo_intensity is 0 then
4:     emo_control, point_control = None, None
5:  else if emo_intensity is -1 then
6:     Use emo_control_pre or default AU values
[…]
8:     Calculate levels $i$ and $j$
9:     Set emo_up_name and emo_down_name
10:     for each AU do
11:        Interpolate emo_control_dict
12:        if mode_type is 3 then
13:           Set point_control_dict
14:        end if
15:     end for
16:  else
17:     Calculate excess $ex$
18:     for each AU do
19:        Extrapolate emo_control_dict
20:        if mode_type is 3 then
21:           Set point_control_dict
22:        end if
23:     end for
24:  end if
D.2 Inference Structure
We present the inference algorithm, which operates in four modes: Mode 0 serves as the base; Mode 1 incorporates audio features to calibrate facial expressions; Mode 2 employs the full method by integrating complete multi-modal features as input; and Mode 3 operates without the AU Source. In Mode 3, to achieve more stable emotion transformations, we employ the point cloud obtained from the Emotion Injector as input to the Isolation Subspace.
Algorithm 2 LES Core Structure
Input: current_mel_input, au, ref, emo_type, emo_intensity
Output: exp_coeff_pred
1:  if au is not None then
2:     Normalize au using means and stds
3:  end if
4:  if mode_type is 0 then
5:     Compute exp_coeff_pred using CrossDimensionAttnNet1 with au and ref
6:  else if mode_type is 1 then
7:     if emo_control is not None then
8:        Adjust au with emo_control
9:     end if
10:     Extract au45 from au
11:     Compute exp_coeff_pred using CrossDimensionAttnNet1 with au and ref
12:     Get pred_ref from PretrainedNetG using current_mel_input, ref, au45
13:     Calculate exp_resi using OffsetDecoder
14:     Add exp_resi to exp_coeff_pred
15:  else if mode_type is 2 then
16:     Compute curr_exp_coeff_pred using CrossDimensionAttnNet1 with au and ref
17:     Calculate emo_resi using CrossDimensionAttnNet2
18:     Calculate exp_resi and audio_embed using OffsetDecoder
19:     Add emo_resi And exp_resi to curr_exp_coeff_pred to get exp_coeff_pred
20:  else if mode_type is 3 then
21:     Extract au45 from au
22:     Get pred_ref from PretrainedNetG using current_mel_input, ref, au45
23:     Calculate exp_resi and audio_embed using OffsetDecoder
24:     Predict aus using AuDecoder
25:     Adjust au45 and combine with aus
26:     Normalize aus
27:     if emo_control is not None then
28:        Adjust aus with emo_control
29:     end if
30:     Compute curr_exp_coeff_pred using CrossDimensionAttnNet1 with aus and ref
31:     if point_control is not None then
32:        Create au_pred
33:        Calculate emo_resi using CrossDimensionAttnNet2
34:        Add emo_resi to curr_exp_coeff_pred
35:     end if
36:     Add exp_resi to curr_exp_coeff_pred to get exp_coeff_pred
37:  end if
38:  return exp_coeff_pred